### v3

@@ 2.2 Proximity Operator and Moreau Envelope
for $\bm{y}^{\star}\in\mathcal{Y}$ to be a stationary point of Problem 1.2.
Theorem 3.2 below presents a relation between (25) and (26). Based on Theorem 3.2, the conditions (25) and (26) are equivalent with $\bm{x}^{\star}=\bm{\varphi}(\bm{y}^{\star})$ under a mild condition (see Corollary 3.4). In a special case where the cost function $F$ is continuously differentiable, a similar relation between the first-order optimality conditions is found in [55, Thm. 2.4]. We note that, even if $C$ is not an embedded submanifold of $\mathcal{X}$ but is a closed Clarke regular subset of $\mathcal{X}$, the same statements in Theorem 3.2 hold.
@@
Consider Problems 1.1 and 1.2. Let $\bm{x}^{\star}\coloneqq\bm{\varphi}(\bm{y}^{\star})\in C$ with some $\bm{y}^{\star}\in\mathcal{Y}$. Then, the following hold:
@@
    Assume $N_{C}(\bm{x}^{\star})\supset\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*})$. If $\bm{y}^{\star}$ satisfies (26), then $\bm{x}^{\star}$ satisfies (25) holds.
@@
Theorem 3.2 and Corollary 3.4 encourage us to focus on finding a stationary point of Problem 1.2 for our goal of finding a stationary point of Problem 1.1. Since Problem 1.2 is an unconstrained optimization problem thanks to the available parametrization $\bm{\varphi}$, Problem 1.2 seems to be more tractable than Problem 1.1 in terms of designing iterative algorithms. Indeed, iterative algorithms with parametrization has been proposed in a case where the cost function $F=h+g\circ\mathfrak{S}$ is smooth [48, 49, 50, 51, 52, 70, 53, 54, 71, 72]. In the next section, we propose an iterative algorithm for Problem 1.2 of nonsmooth cost function $F$ and parametrization $\bm{\varphi}$.
4 Variable Smoothing Algorithm for Problem 1.2
4.1 Properties of smoothed surrogate function for Problem 1.2
In this section, we investigate properties of our smoothed surrogate function $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}\ (\mu\in(0,\eta^{-1}))$ of the cost function $(h+g\circ\mathfrak{S})\circ\bm{\varphi}$ in Problem 1.2, where $^{\mu}g$ is the Moreau envelope of $g$ in (21). The goal of this section is to present an asymptotic property, called gradient (sub-)consistency, of the gradient of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}\ (\mu\searrow 0)$ to approximate subgradients of $(h+g\circ\mathfrak{S})\circ\bm{\varphi}$ (see Theorem 4.4 and Remark 4.5). By exploiting the gradient (sub-)consistency of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$, we will establish an asymptotic convergence analysis for the proposed variable smoothing algorithm in Section 4.2. We summarize basic properties of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$, with $\mu\in(0,\eta^{-1})$ fixed, which follow from direct applications of Fact 2.7 (d).
@@
For a smoothed surrogate function $F^{\langle\mu\rangle}\circ\bm{\varphi}\coloneqq(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}\ (\mu\in(0,\eta^{-1}))$ of $F\circ\bm{\varphi}=(h+g\circ\mathfrak{S})\circ\bm{\varphi}$ in Problem 1.2, we have
@@
$(\mu\in(0,\eta^{-1}))\quad$ $-\infty<\inf_{\bm{y}\in\mathcal{Y}}F^{\langle\mu\rangle}\circ\bm{\varphi}(\bm{y})$ (29)
with a Lipschitz constant $L_{g}>0$ of $g$. Moreover, with $0<\mu_{2}<\mu_{1}<\eta^{-1}$, we have
@@
The inequality (28) follows from (23). The inequality (29) is verified by (28) and $\inf_{\bm{y}\in\mathcal{Y}}F\circ\bm{\varphi}(\bm{y})>$-$\infty$ from the condition (iv) in Problem 1.1. It suffices for deriving (30) to show
@@
The Lipschitz continuity of the gradient of a cost function is one of important ingredients for optimization to ensure the descent lemma (see, e.g., [73, Lemma 5.7]). The following presents a sufficient condition for the Lipschitz constant of the gradient of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ (see also Remark 4.3 for its further sufficient conditions).
@@
Consider a smoothed surrogate function $F^{\langle\mu\rangle}\circ\bm{\varphi}\coloneqq(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}\ (\mu\in(0,\eta^{-1}))$ of $F\circ\bm{\varphi}=(h+g\circ\mathfrak{S})\circ\bm{\varphi}$ in Problem 1.2, and the following conditions:
@@
    (Sufficient conditions regarding $\bm{\varphi}$) The conditions (A4) and (A5) are automatically satisfied in a case where $C$ is a linear subspace of $\mathcal{X}$ and $\bm{\varphi}$ is a linear operator. Even for the other cases, we will present an example of $(C,\bm{\varphi})$, in Corollary 5.1 (see also Lemma D.2), achieving the conditions (A4) and (A5).
From (27), an approximation of a stationary point of Problem 1.2 can be reformulated as finding $\bm{y}^{\star}\in\mathcal{Y}$ where the value $d\left(\bm{0},\partial_{\rm L}(F\circ\bm{\varphi})(\bm{y}^{\star})\right)$ with $F=h+g\circ\mathfrak{S}$ is sufficiently small. Theorem 4.4 below presents an upper bound of this distance with the gradient of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ in (37) by showing the gradient sub-consistency of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ for $F$ (see also Remark 4.5).
@@
    Consider a smoothed surrogate function $F^{\langle\mu\rangle}\circ\bm{\varphi}\coloneqq(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}\ (\mu\in(0,\eta^{-1}))$ of $F\circ\bm{\varphi}=(h+g\circ\mathfrak{S})\circ\bm{\varphi}$ in Problem 1.2. Then, we have
@@
    Although the converse inclusion of (35) is not required for this paper, the equality
@@
    called the gradient consistency property [35] of $^{\mu}J\circ\mathcal{F}$ for $J\circ\mathcal{F}$, can also be shown (i) by checking that $^{\mu}J\circ\mathcal{F}$ epi-converges to $J\circ\mathcal{F}$ (see [57, Def. 7.1] for the definition of epi-convergence); and (ii) by applying [61, Lemma 3.4] to $^{\mu}J\circ\mathcal{F}$. Indeed, (i) can be verified by [57, Thm. 7.4(d)] and $^{\mu}J\circ\mathcal{F}\nearrow J\circ\mathcal{F}$ as $\mu\searrow 0$.
@@
    In a special case where $J$ is convex, we can check that $^{\mu}J\circ\mathcal{F}$ in Theorem 4.4 (a) has the gradient consistency property for $J\circ\mathcal{F}$ by using [61, Cor. 5.8]. However, the discussion in the proof of [61, Cor. 5.8] can not apply to our case where $J$ is just weakly convex because the proof of [61, Cor. 5.8] relies on the structure of the convexity of $J$ (see Attouch’s theorem [57, Thm. 12.35]). In order to show the gradient consistency of $^{\mu}J\circ\mathcal{F}$ even with $J$ weakly convex, we take a different approach in the proof of Theorem 4.4.
@@ 4.2 Variable Smoothing Algorithm with Convergence Analysis
The inequality (37) in Theorem 4.4 (b) implies that finding $\bm{y}^{\star}\in\mathcal{Y}$ satisfying (27) can be reduced to
@@
In this section, for Problem 1.2 via the problem in (39), we present a variable smoothing algorithm in Algorithm 1 as an extension of the variable smoothing algorithm [23] that has been proposed for Problem 1.1 with a linear operator $\mathfrak{S}$ and $C\coloneqq\mathcal{X}$. The proposed algorithm is designed with a time-varying gradient descent update for minimization of
@@
at $n$th update from $\bm{y}_{n}\in\mathcal{Y}$ to $\bm{y}_{n+1}\coloneqq\bm{y}_{n}-\gamma_{n}\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\in\mathcal{Y}$ with $\mu_{n}\searrow 0$ and a stepsize $\gamma_{n}>0$. In a case where $\mathrm{prox}_{\mu_{n}g}$ has a closed-form expression, the values of $F_{n}\circ\bm{\varphi}(\widebar{\bm{y}})$ and $\nabla(F_{n}\circ\bm{\varphi})(\widebar{\bm{y}})$ are available for every $\widebar{\bm{y}}\in\mathcal{Y}$ (see (21) and Fact 2.7 (b)).
@@
where $\eta>0$ is the level of the weak convexity of $g$ (see Problem 1.1). We have many choices of $(\mu_{n})_{n=1}^{\infty}$ enjoying (41) as shown in Example 4.6 (see Remark 4.10 for a reasonable choice of $(\mu_{n})_{n=1}^{\infty}$).
Algorithm 1 Variable smoothing for Problem 1.2 (VSmooth)
@@
3:while stopping criteria are not satisfied do
@@
5:  Find $\gamma_{n}>0$ satisfying (48) with $c$ (see Example 4.8)
6:  $\bm{y}_{n+1}\leftarrow\bm{y}_{n}-\gamma_{n}\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})$
@@
Every stepsize $\gamma_{n}>0$ in Algorithm 1 is chosen to enjoy the so-called Armijo condition (see, e.g., [74, Thm. 1.1]) with a predetermined constant $c\in(0,1)$:
@@
Consider Problem 1.2 and Alg. 1. For $F_{n}=h+{}^{\mu_{n}}g\circ\mathfrak{S}$, we assume:
@@
Under Assumption 4.7 (a), we have mainly two choices of $\gamma_{n}$ achieving (48) (see, e.g., [74, Thm. 1.1]):
@@
    The so-called backtracking algorithm in Algorithm 2 returns $\gamma_{n}$ enjoying the Armijo condition (45) with $c\in(0,1)$, where $\gamma_{\rm initial}\in\mathbb{R}_{++}$ is an initial guess for $\gamma_{n}$, and $\rho\in(0,1)$. In this case, $\gamma_{n}$ is known to have the following lower bound:
    $(n\in\mathbb{N})\quad\gamma_{n}\geq\min\left\{\gamma_{\rm initial}L_{\nabla(F_{n}\circ\bm{\varphi})},2\rho(1-c)\right\}L_{\nabla(F_{n}\circ\bm{\varphi})}^{-1}\geq\beta L_{\nabla(F_{n}\circ\bm{\varphi})}^{-1}$ (50)
    with $\beta\coloneqq\min\left\{\gamma_{\rm initial}L_{\nabla(F_{1}\circ\bm{\varphi})},2\rho(1-c)\right\}\in\mathbb{R}_{++}$, where the first inequality follows by [74, Thm. 1.1], and the second inequality follows [^9] The derivation of the second inequality is inspired by [24, the equality just after (4.12)]. by $L_{\nabla(F_{n}\circ\bm{\varphi})}\geq L_{\nabla(F_{1}\circ\bm{\varphi})}$ (see the sentence just after (47)). From (50), Algorithm 2 is guaranteed to terminate in at most $\max\left\{1,\left\lceil\log_{\rho}\left(\frac{2(1-c)}{\gamma_{\rm inital}L_{\nabla(F_{n}\circ\bm{\varphi})}}\right)\right\rceil\right\}$ updates, and $\gamma_{n}$ given by Algorithm 2 satisfies (48), where $\lceil\cdot\rceil$ stands for the ceiling function. We note that any knowledge on $L_{\nabla(F_{n}\circ\bm{\varphi})}$ is not required in Algorithm 2.
@@
$c\in(0,1),\rho\in(0,1),\ \gamma_{\rm initial}\in\mathbb{R}_{++}$
$\gamma_{n}\leftarrow\gamma_{\rm initial}$
while $F_{n}\circ\bm{\varphi}(\bm{y}_{n}-\gamma_{n}\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n}))>F_{n}\circ\bm{\varphi}(\bm{y}_{n})-c\gamma_{n}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert^{2}$ do
@@
Theorem 4.9 (Convergence analysis of Algorithm 1).
Consider Problem 1.2. Choose arbitrarily $\bm{y}_{1}\in\mathcal{Y}$, $c\in(0,1)$, and $(\mu_{n})_{n=1}^{\infty}\subset(0,(2\eta)^{-1}]$ satisfying (41). Suppose that $(\bm{y}_{n})_{n=1}^{\infty}\subset\mathcal{Y}$ is generated by Algorithm 1 under Assumption 4.7 with $F_{n}\coloneqq F^{\langle\mu_{n}\rangle}\coloneqq h+{}^{\mu_{n}}g\circ\mathfrak{S}\ (n\in\mathbb{N})$. Then, $F_{n}\circ\bm{\varphi}\ (n\in\mathbb{N})$ and $(\bm{y}_{n})_{n=1}^{\infty}$ enjoy the following:

### v4

@@ 2.2 Proximity Operator and Moreau Envelope
for $\bm{y}^{\star}\in\mathcal{Y}$ to be a stationary point of Problem 1.4.
Theorem 3.2 below presents a relation between (25) and (26). Based on Theorem 3.2, the conditions (25) and (26) are equivalent with $\bm{x}^{\star}=\bm{\varphi}(\bm{y}^{\star})$ under a certain condition (see Corollary 3.4). In a special case where the cost function $F$ is continuously differentiable, a similar relation between the first-order optimality conditions is found in [56, Thm. 2.4]. We note that, even if $C$ is not an embedded submanifold of $\mathcal{X}$ but is just a Clarke regular subset of $\mathcal{X}$, the same statements in Theorem 3.2 hold.
@@
Consider Problems 1.1 and 1.4. Let $\bm{x}^{\star}\coloneqq\bm{\varphi}(\bm{y}^{\star})\in C$ with some $\bm{y}^{\star}\in\mathcal{Y}$. Then, the following hold:
@@
    Assume $N_{C}(\bm{x}^{\star})\supset\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*})$. If $\bm{y}^{\star}$ satisfies (26), then $\bm{x}^{\star}$ satisfies (25).
@@
Remark 3.6 (On submersion condition).
A parametrization $\bm{\varphi}$ of $C$ in Problem 1.1 does not automatically satisfy the submersion condition $\mathrm{Ran}(\mathrm{D}\bm{\varphi}(\bm{y}))=T_{C}(\bm{\varphi}(\bm{y}))$ at each point $\bm{y}\in\mathcal{Y}$ (see, e.g., [47] and footnote21 in Appendix A.2, for a parametrization with the submersion condition). However, even in the absence of the submersion condition [^9] For a manifold $C$ in Problem 1.1, there exists a parametrization satisfying the submersion condition locally. Indeed, for $\bm{x}\in C$, the exponential map $\mathrm{Exp}_{\bm{x}}:T_{C}(\bm{x})\to C$ (see Example 1.3 (c)) is a local diffeomorphism at $\bm{0}\in T_{C}(\bm{x})$ [55, Lemma 6.16], hence submersion over an open neighborhood of $\bm{0}$. By using a family of parametrizations satisfying the submersion condition locally, an adaptive parametrization strategy has also been proposed [58, 29, 50] for the smooth case of the cost function. of $\bm{\varphi}$, Theorem 3.2 (a) ensures that the first-order optimality condition (26) of Problem 1.4 at $\bm{y}^{\star}\in\mathcal{Y}$ is a necessary condition for the first-order optimality condition (25) of Problem 1.1 at $\bm{x}^{\star}\coloneqq\bm{\varphi}(\bm{y}^{\star})\in C$, hence is a necessary condition for the local optimality of Problem 1.1 at $\bm{x}^{\star}$.
Theorem 3.2 and Corollary 3.4 encourage us to focus on finding a stationary point of Problem 1.4 for our goal of finding a stationary point of Problem 1.1. Since Problem 1.4 is an unconstrained optimization problem thanks to the parametrization $\bm{\varphi}$, Problem 1.4 seems to be more tractable than Problem 1.1 in terms of designing iterative algorithms. Indeed, iterative algorithms with parametrization have been proposed in a case where the cost function $F=h+g\circ\mathfrak{S}$ is smooth [84, 36, 43, 41, 58, 39, 47, 48, 29, 50]. In Section 4, along such parametrization strategies, we propose an iterative algorithm applicable to Problem 1.4 even with a nonsmooth cost function $F$.
We remark that, as an alternative strategy to the parametrization strategy for translating Problem 1.1 into an unconstrained problem, penalty-based strategies have been proposed, e.g., [65, Ch. 12][82], in a case where $C$ is defined as equality constraints. See Appendix A.2 for a further discussion of the penalty-based and parametrization strategies.
4 Variable Smoothing Algorithm for Problem 1.4
4.1 Properties of smoothed surrogate function for Problem 1.4
In this section, we investigate properties of our smoothed surrogate function $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}\ (\mu\in(0,\eta^{-1}))$ of the cost function $(h+g\circ\mathfrak{S})\circ\bm{\varphi}$ in Problem 1.4, where $^{\mu}g$ is the Moreau envelope of $g$ in (21). The goal of this section is to present an asymptotic property, called gradient (sub-)consistency, of the gradient of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}\ (\mu\searrow 0)$ to approximate subgradients of $(h+g\circ\mathfrak{S})\circ\bm{\varphi}$ (see Theorem 4.4 and Remark 4.5). By exploiting the gradient (sub-)consistency of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$, we will establish an asymptotic convergence analysis for the proposed variable smoothing algorithm in Section 4.2. We summarize basic properties of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$, with $\mu\in(0,\eta^{-1})$ fixed, which follow from direct applications of Fact 2.7 (d).
@@
Let $F^{\langle\mu\rangle}\circ\bm{\varphi}\coloneqq(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}\ (\mu\in(0,\eta^{-1}))$ be a smoothed surrogate function of $F\circ\bm{\varphi}=(h+g\circ\mathfrak{S})\circ\bm{\varphi}$ in Problem 1.4. Then, for $\widebar{\bm{y}}\in\mathcal{Y}$ and $\mu\in(0,\eta^{-1})$, we have the following expressions by (21) and Fact 2.7 (b):
$\begin{cases}F^{\langle\mu\rangle}\circ\bm{\varphi}(\widebar{\bm{y}})=h(\bm{\varphi}(\widebar{\bm{y}}))+g\left(\mathrm{prox}_{\mu g}\left(\mathfrak{S}(\bm{\varphi}(\widebar{\bm{y}}))\right)\right)+\frac{1}{2\mu}\left\lVert\mathrm{prox}_{\mu g}\left(\mathfrak{S}\left(\bm{\varphi}(\widebar{\bm{y}})\right)\right)-\mathfrak{S}\left(\bm{\varphi}(\widebar{\bm{y}})\right)\right\rVert^{2},\\ \nabla(F^{\langle\mu\rangle}\circ\bm{\varphi})(\widebar{\bm{y}})=(\mathrm{D}\bm{\varphi}(\bm{y}))^{*}\left[\nabla h(\bm{\varphi}(\widebar{\bm{y}}))+\left(\mathrm{D}\mathfrak{S}\left(\bm{\varphi}(\widebar{\bm{y}})\right)\right)^{*}\left[\frac{\mathfrak{S}(\bm{\varphi}(\widebar{\bm{y}}))-\mathrm{prox}_{\mu g}\left(\mathfrak{S}(\bm{\varphi}(\widebar{\bm{y}}))\right)}{\mu}\right]\right].\end{cases}$ (28)
With a Lipschitz constant $L_{g}>0$ of $g$, we have
@@
$(\mu\in(0,\eta^{-1}))\quad$ $-\infty<\inf_{\bm{y}\in\mathcal{Y}}F^{\langle\mu\rangle}\circ\bm{\varphi}(\bm{y}).$ (30)
Moreover, with $0<\mu_{2}<\mu_{1}<\eta^{-1}$, we have
@@
The expressions (28) follow from (21) and Fact 2.7 (b). The inequality (29) follows from (23). The inequality (30) is verified by (29) and $\inf_{\bm{y}\in\mathcal{Y}}F\circ\bm{\varphi}(\bm{y})>$-$\infty$ from the condition (iv) in Problem 1.1. To derive (31), it suffices to show
@@
The Lipschitz continuity of the gradient of a cost function is one of important ingredients for optimization to ensure the descent lemma (see, e.g., [9, Lemma 5.7]). The following presents a sufficient condition for the Lipschitz constant of the gradient of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ (see also Remark 4.3 for its further sufficient conditions).
@@
Consider a smoothed surrogate function $F^{\langle\mu\rangle}\circ\bm{\varphi}\coloneqq(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}\ (\mu\in(0,\eta^{-1}))$ of $F\circ\bm{\varphi}=(h+g\circ\mathfrak{S})\circ\bm{\varphi}$ in Problem 1.4, and the following conditions:
@@
    (Sufficient conditions regarding $\bm{\varphi}$) The conditions (A4) and (A5) are automatically satisfied in a case where $C$ is a linear subspace of $\mathcal{X}$ and $\bm{\varphi}$ is a linear operator. Even for the other cases, we will present an example of $(C,\bm{\varphi})$, in Corollary 5.1 (see also Lemma E.2), achieving the conditions (A4) and (A5).
From (27) and Theorem 3.2, we aim to find a stationary point $\bm{y}^{\star}\in\mathcal{Y}$ of Problem 1.4 such that $d\left(\bm{0},\partial_{\rm L}(F\circ\bm{\varphi})(\bm{y}^{\star})\right)=0$ with $F=h+g\circ\mathfrak{S}$. Theorem 4.4 below presents an upper bound of this distance with the gradient of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ in (38) by showing the gradient sub-consistency of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ for $F$ (see also Remark 4.5).
@@
    Consider a smoothed surrogate function $F^{\langle\mu\rangle}\circ\bm{\varphi}\coloneqq(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}\ (\mu\in(0,\eta^{-1}))$ of $F\circ\bm{\varphi}=(h+g\circ\mathfrak{S})\circ\bm{\varphi}$ in Problem 1.4 (see also Problem 1.1). Then, we have
@@
    Although the converse inclusion of (36) is not required in this paper, the equality
@@
    called the gradient consistency property [26] of $^{\mu}J\circ\mathcal{F}$ for $J\circ\mathcal{F}$, can also be shown (i) by checking that $^{\mu}J\circ\mathcal{F}$ epi-converges to $J\circ\mathcal{F}$ (see [69, Def. 7.1] for the definition of epi-convergence); and (ii) by applying [18, Lemma 3.4] to $^{\mu}J\circ\mathcal{F}$. Indeed, (i) can be verified by [69, Thm. 7.4(d)] and $^{\mu}J\circ\mathcal{F}\nearrow J\circ\mathcal{F}$ as $\mu\searrow 0$.
@@
    In a special case where $J$ is convex, we can check that $^{\mu}J\circ\mathcal{F}$ in Theorem 4.4 (a) has the gradient consistency property for $J\circ\mathcal{F}$ by using [18, Cor. 5.8]. However, the discussion in the proof of [18, Cor. 5.8] cannot apply to our case where $J$ is just weakly convex because the proof of [18, Cor. 5.8] relies on the structure of the convexity of $J$ (see Attouch’s theorem [69, Thm. 12.35]). In order to show the gradient consistency of $^{\mu}J\circ\mathcal{F}$ even with $J$ weakly convex, we take a different approach in the proof of Theorem 4.4. We note that our smoothed surrogate function $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ with a linear $\mathfrak{S}$ and $\bm{\varphi}\coloneqq\mathrm{Id}$ has been employed in the existing variable smoothing algorithms [13, 10, 66], although the gradient consistency property has not been discussed therein.
  3. (c)
    We can construct another smoothed surrogate function, with gradient consistency property, of $F\circ\bm{\varphi}$ via the integral convolution of $F\circ\bm{\varphi}$ and a kernel function, called mollifier (see [69, Exm. 7.19, Thm. 9.67] or [19, Thm. 4.7]), because $F\circ\bm{\varphi}$ is subdifferentially regular (see Lemma 2.4). However, it is questionable to obtain closed-form expressions of this type of smoothed surrogate functions for $F\circ\bm{\varphi}$ and its gradient, except for simple functions, e.g., the plus function $\mathbb{R}\ni x\mapsto\max\{x,0\}$ [22]. In contrast, the closed-form expressions of our smoothed surrogate function $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ and its gradient are available as computable tools in a case where $g$ is prox-friendly (see (28)).
@@ 4.2 Variable Smoothing Algorithm with Convergence Analysis
The inequality (38) in Theorem 4.4 (b) implies that finding a stationary point $\bm{y}^{\star}\in\mathcal{Y}$ of Problem 1.4 satisfying (27) can be reduced to the following problem:
@@
In this section, for finding a stationary point of Problem 1.4 via the problem in (40), we present a variable smoothing algorithm in Algorithm 1 as an extension of the variable smoothing algorithm [13] that has been proposed for Problem 1.4 with $\bm{\varphi}=\mathrm{Id}$ and a linear operator $\mathfrak{S}$. We note that the algorithm [13] is not applicable to Problem 1.4 due to the nonlinearity of $\mathfrak{S}$ even if $\bm{\varphi}=\mathrm{Id}$. The proposed algorithm is designed with a time-varying gradient descent update for minimization of
@@
at $n$th update from $\bm{y}_{n}\in\mathcal{Y}$ to $\bm{y}_{n+1}\coloneqq\bm{y}_{n}-\gamma_{n}\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\in\mathcal{Y}$ with $\mu_{n}\searrow 0$ and a stepsize $\gamma_{n}>0$. In a case where $\mathrm{prox}_{\mu_{n}g}$ has a closed-form expression, the values of $F_{n}\circ\bm{\varphi}(\widebar{\bm{y}})$ and $\nabla(F_{n}\circ\bm{\varphi})(\widebar{\bm{y}})$ are available for every $\widebar{\bm{y}}\in\mathcal{Y}$ (see (28)).
@@
where $\eta>0$ is the level of the weak convexity of $g$ (see Problem 1.1). We have many choices of $(\mu_{n})_{n=1}^{\infty}$ enjoying (42) as shown in Example 4.6.
Algorithm 1 Variable smoothing for Problem 1.4 (VSmooth)
@@
3:while stopping criterion in not satisfied do
@@
5:  Set $\gamma_{n}>0$ satisfying (49) with $c$ (see Example 4.8)
6:  $\bm{y}_{n+1}\leftarrow\bm{y}_{n}-\gamma_{n}\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})$ (see (28) for the expression of $\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})$)
@@
Every stepsize $\gamma_{n}>0$ in Algorithm 1 is chosen to enjoy the so-called Armijo condition (see, e.g., [5, Thm. 1.1]) with a predetermined constant $c\in(0,1)$:
@@
Consider Problem 1.4 and Alg. 1. For $F_{n}=h+{}^{\mu_{n}}g\circ\mathfrak{S}$, we assume:
@@
Under Assumption 4.7 (a), we have mainly two choices of $\gamma_{n}$ achieving (49) (see, e.g., [5, Thm. 1.1]):
@@
    The so-called backtracking algorithm in Algorithm 2 returns $\gamma_{n}$ enjoying the Armijo condition (46) with $c\in(0,1)$, where $\gamma_{\rm init}\in\mathbb{R}_{++}$ is an initial guess for $\gamma_{n}$, and $\rho\in(0,1)$. In this case, $\gamma_{n}$ is known to have the following lower bound:
    $(n\in\mathbb{N})\quad\gamma_{n}\geq\min\left\{\gamma_{\rm init}L_{\nabla(F_{n}\circ\bm{\varphi})},2\rho(1-c)\right\}L_{\nabla(F_{n}\circ\bm{\varphi})}^{-1}\geq\beta L_{\nabla(F_{n}\circ\bm{\varphi})}^{-1}$ (51)
    with $\beta\coloneqq\min\left\{\gamma_{\rm init}L_{\nabla(F_{1}\circ\bm{\varphi})},2\rho(1-c)\right\}\in\mathbb{R}_{++}$, where the first inequality follows by [5, Thm. 1.1], and the second inequality follows [^10] The derivation of the second inequality is inspired by [10, the equality just after (4.12)]. by $L_{\nabla(F_{n}\circ\bm{\varphi})}\geq L_{\nabla(F_{1}\circ\bm{\varphi})}$ (see the sentence just after (48)). From (51), Algorithm 2 is guaranteed to terminate in at most $\max\left\{1,\left\lceil\log_{\rho}\left(\frac{2(1-c)}{\gamma_{\rm init}L_{\nabla(F_{n}\circ\bm{\varphi})}}\right)\right\rceil\right\}=\max\left\{1,\left\lceil\log_{\rho}\left(\frac{2\mu_{n}(1-c)}{\gamma_{\rm init}(\mu_{n}\varpi_{1}+\varpi_{2})}\right)\right\rceil\right\}$ updates, and $\gamma_{n}$ given by Algorithm 2 satisfies (49), where $\lceil\cdot\rceil$ stands for the ceiling function. We note that any knowledge on $L_{\nabla(F_{n}\circ\bm{\varphi})}=\varpi_{1}+\varpi_{2}\mu_{n}^{-1}$ with $\varpi_{1},\varpi_{2}\in\mathbb{R}_{+}$ is not required in Algorithm 2.
@@
$c\in(0,1),\rho\in(0,1),\ \gamma_{\rm init}\in\mathbb{R}_{++}$
$\gamma_{n}\leftarrow\gamma_{\rm init}$
$f_{n}\leftarrow F_{n}\circ\bm{\varphi}(\bm{y}_{n})\in\mathbb{R}$, $\bm{d}_{n}\leftarrow\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\in\mathcal{Y}$
while $F_{n}\circ\bm{\varphi}(\bm{y}_{n}-\gamma_{n}\bm{d}_{n})>f_{n}-c\gamma_{n}\left\lVert\bm{d}_{n}\right\rVert^{2}$ do
@@
Theorem 4.9 (Asymptotic convergence analysis of Algorithm 1).
Consider Problem 1.4. Choose arbitrarily $\bm{y}_{1}\in\mathcal{Y}$, $c\in(0,1)$, and $(\mu_{n})_{n=1}^{\infty}\subset(0,(2\eta)^{-1}]$ satisfying (42). Suppose that $(\bm{y}_{n})_{n=1}^{\infty}\subset\mathcal{Y}$ is generated by Algorithm 1 under Assumption 4.7 with $F_{n}\coloneqq F^{\langle\mu_{n}\rangle}\coloneqq h+{}^{\mu_{n}}g\circ\mathfrak{S}\ (n\in\mathbb{N})$. Then, $F_{n}\circ\bm{\varphi}\ (n\in\mathbb{N})$ and $(\bm{y}_{n})_{n=1}^{\infty}$ enjoy the following: